OpenStudy (anonymous):

Find the missing terms and write them in decreasing order of exponents. 21m7 − 7m5 + 49m = 7m(

OpenStudy (danjs):

each one of those contains a common factor of 7 and m factoring is similar to doing this, 7m/7m = 1, multiplying by 1, nothing changed \[\large \frac{ 7m }{ 7m }*(21m^7-7m^5+49m)\]

OpenStudy (danjs):

notice each of those terms can be divided by 7m cleanly \[\large 7m*(\frac{ 21m^7 }{ 7m }-\frac{ 7m^5 }{ 7m }+\frac{ 49m }{ 7m })\]

OpenStudy (anonymous):

Yes, I understand thank you so much!

OpenStudy (danjs):

giving \[\large 7m * (3m^6 - 1m^4+7)\]

OpenStudy (danjs):

k, welcome